OpenStudy (anonymous):

solve y=x^3 for y=mxtb Does this even have a solution???

jimthompson5910 (jim_thompson5910):

you mean put y=x^3 into y = mx+b form?

OpenStudy (anonymous):

ya lol

jimthompson5910 (jim_thompson5910):

you can't do it

jimthompson5910 (jim_thompson5910):

it's impossible to write y = x^3 in y = mx+b form

OpenStudy (anonymous):

thnx

jimthompson5910 (jim_thompson5910):

simply because y = x^3 is a cubic equation and y = mx+b is a linear equation, they're two completely different types of equations

OpenStudy (anonymous):

lol ive learned more than i did at school here at home

jimthompson5910 (jim_thompson5910):

lol that's good

OpenStudy (anonymous):

thnx